OpenStudy (erinweeks):

One more question tonight! A statement Sn about the positive integers is given. Write statements Sk and Sk+1, simplifying Sk+1 completely. Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3

ganeshie8 (ganeshie8):

Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3 Sk : ?

ganeshie8 (ganeshie8):

write the same statemetn as-it-is, just replace n wid k,

OpenStudy (erinweeks):

i dont know what the hell sk is?

OpenStudy (erinweeks):

okay but than how can i solve it with k?

ganeshie8 (ganeshie8):

we dont need to solve it. read the question once, it is asking us to JUST WRITE the statement for Sk and Sk+1

ganeshie8 (ganeshie8):

Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3 Sk : 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + k(k + 1) = [k(k + 1)(k + 2)]/3

ganeshie8 (ganeshie8):

thats it, we're done wid Sk work Sk+1 now

OpenStudy (erinweeks):

how do i work sk+1 tho?

ganeshie8 (ganeshie8):

put k+1 in place of n

ganeshie8 (ganeshie8):

Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3 Sk+1 : ?

ganeshie8 (ganeshie8):

give it a try

OpenStudy (erinweeks):

so its going to be sk+1(sk+1 + 1) = {sk+1(sk+1 +1)(sk+1 + 2)] / 3

ganeshie8 (ganeshie8):

dont put s

OpenStudy (erinweeks):

its just 1?

ganeshie8 (ganeshie8):

and where are 1.2+ 2.3 + 3.4.... in the start? write it properly/completely

OpenStudy (erinweeks):

sorry :(

ganeshie8 (ganeshie8):

its okay... try if u can :)

ganeshie8 (ganeshie8):

See how we wrote the statemetn Sk above, Sk+1 should look something like that

OpenStudy (erinweeks):

is it just 1 tho i put in?

OpenStudy (erinweeks):

k+1?

ganeshie8 (ganeshie8):

yes replace n wid k+1

OpenStudy (erinweeks):

1*2+2*3+3*4+....+ k+1(k+1 + 1) = [k+1(k+1+1)(k+1 +2)] / 3

ganeshie8 (ganeshie8):

\(S\color{red}{n}: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + \color{red}{n}(\color{red}{n} + 1) = [\color{red}{n}(\color{red}{n} + 1)(\color{red}{n} + 2)]/3\) \(S\color{red}{k+1} : 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + (\color{red}{k+1})(\color{red}{k+1} + 1) = [(\color{red}{k+1})(\color{red}{k+1} + 1)(\color{red}{k+1} + 2)]/3\)

ganeshie8 (ganeshie8):

Yup ! sumplify

OpenStudy (erinweeks):

i have to simplify -____-

ganeshie8 (ganeshie8):

Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3 Sk+1 : 1*2+2*3+3*4+....+ (k+1)(k+2) = [(k+1)(k+2)(k+3)] / 3

OpenStudy (erinweeks):

flutter my life -___-

OpenStudy (erinweeks):

f*ck

ganeshie8 (ganeshie8):

:|

OpenStudy (erinweeks):

thats it?

ganeshie8 (ganeshie8):

nope

OpenStudy (erinweeks):

f*ckkkk

ganeshie8 (ganeshie8):

Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3 Sk+1 : 1*2+2*3+3*4+....+ (k+1)(k+2) = [(k+1)(k+2)(k+3)] / 3 ^^^^^^^^^^^^^^

ganeshie8 (ganeshie8):

simplify that, then u may deserve some good sleep :)

OpenStudy (erinweeks):

[(k+4)]

ganeshie8 (ganeshie8):

put it in wolfram

OpenStudy (erinweeks):

i dont know how to use that

ganeshie8 (ganeshie8):

this expression :- (k+1)(k+2)(k+3)

ganeshie8 (ganeshie8):

open wolframalpha

OpenStudy (erinweeks):

i did but is it the alternate forms???

OpenStudy (erinweeks):

yea i got the same thing but which one is the answer lol it confuses me

OpenStudy (erinweeks):

k+6?

ganeshie8 (ganeshie8):

Yup ! its in Alternate forms : 2nd line (k+1)(k+2)(k+3) = k^3+6k^2+11k+6

OpenStudy (erinweeks):

okay and thats simplified form.. so i put (k+1)(k+2)= k^3+6k^2+11k+6

ganeshie8 (ganeshie8):

Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3 Sk+1 : 1*2+2*3+3*4+....+ (k+1)(k+2) = [(k+1)(k+2)(k+3)] / 3 : 1*2+2*3+3*4+....+ (k+1)(k+2) = [k^3+6k^2+11k+6] / 3

ganeshie8 (ganeshie8):

thats right ! ive put the same above :)

ganeshie8 (ganeshie8):

u may divide by 3 if u want, if u aren 't sleepy lol

OpenStudy (erinweeks):

nah thats too much im half asleep typing lol ..

ganeshie8 (ganeshie8):

Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3 Sk+1 : 1*2+2*3+3*4+....+ (k+1)(k+2) = [(k+1)(k+2)(k+3)] / 3 : 1*2+2*3+3*4+....+ (k+1)(k+2) = [k^3+6k^2+11k+6] / 3 : 1*2+2*3+3*4+....+ (k+1)(k+2) = k^3/3+2k^2+11k/3+2

ganeshie8 (ganeshie8):

thats completely simplified... have good sleep :)

OpenStudy (erinweeks):

thank you soo muchh (:

ganeshie8 (ganeshie8):

np :D